OpenStudy (anonymous):

Convert 143 over 7 to a mixed number

OpenStudy (anonymous):

20 3/7

hero (hero):

\[\frac{143}{7} = \frac{140 + 3}{7} = \frac{140}{7} + \frac{3}{7} = 20 + \frac{3}{7}\]

OpenStudy (anonymous):

Thanks you guy!

hero (hero):

yw

OpenStudy (anonymous):

Yw :)